# Continual Learning in Medical Imaging from Theory to Practice: A Survey and Practical Analysis

### v1

@@ Abstract
keywords:
[…]
1 Introduction
@@
Deep Learning (DL) algorithms are rapidly gaining relevance in medical imaging, enabling segmentation (Chen et al., 2020; Hofmanninger et al., 2020b), classification (Kumar et al., 2023), and detection Halder et al. (2020) of anatomical structures and anomalies (Schlegl et al., 2019; Qazi et al., 2023) relevant for diagnosis, prediction, or prognosis. In some cases, their capabilities surpass those of human experts (Walsh et al., 2018; De Fauw et al., 2018), making them a central tool in the advancement of the use of medical data to support clinical decisions (Ahsan et al., 2022; Adlung et al., 2021). Although there are several DL works in medicine, the task is often oversimplified by assuming a static, centralized data source and neglecting the dynamic nature of the evolving data.
In general, a model is trained and tested on different splits of the same dataset, and the models are expected to generalize well to other datasets of the same nature. However, this does not reflect the reality of real-world deployment because of frequent changes in data distribution across various institutions, diagnostic devices, and population demographics (Cohen et al., 2021). For Instance, in 2020, Google developed a DL system trained on a vast heterogeneous dataset that achieved a high performance in an offline setting for detecting diabetic eye diseases (Beede et al., 2020). However, when this system is deployed, its performance drops drastically, highlighting the severity of the problem and emphasizing the need for robust systems. These issues can be partially addressed through domain adaptation and generalization.
@@
Continual learning (CL) has emerged as a promising approach for training deep neural networks in order to overcome these challenges. CL offers a solution that enables neural networks to learn continuously, resulting in a more resilient model capable of accommodating new tasks without forgetting previous tasks. Although the model remains static when deployed in a static environment, it continuously updates itself in a CL setting, thereby yielding a more robust and versatile model. Addressing the challenges mentioned above requires consideration of the rigidity and plasticity of the neural networks. Rigidity, the ability of a neural network to preserve previous knowledge, and plasticity, the ability to learn new tasks are crucial factors. A model with high plasticity can learn new tasks over time, achieving good performance on future tasks but deteriorating performance on older tasks, a phenomenon known as "catastrophic forgetting". An ideal neural network should balance rigidity and plasticity to demonstrate its ability to learn new tasks without forgetting prior knowledge. CL has emerged as a promising approach for balancing these properties, enabling neural networks to learn continuously and adapt to evolving data distribution.
1.1 Motivation
@@
Significant efforts have been made to develop CL techniques for the natural and medical imaging domains (Douillard et al., 2022; Wang et al., 2022b, a; Smith et al., 2023; Qazi et al., 2024). However, numerous unique challenges arise when dealing with medical data. As illustrated in Figure 1, medical imaging often encounters the challenge of changes in data distribution over time owing to (a) the incorporation of data from new medical centers that might vary in acquisition equipment and patient demographics and (b) modifications that could impact the equipment available at the same site, such as the degradation of current equipment or replacement of devices. (c) Integrating new modality streams over time could be advantageous for downstream tasks. (d) The addition of new tasks may also prompt alterations. Several studies have been conducted to address these challenges. In this work, we aim to conduct a survey of the existing literature and research that may highlight the solutions, the practicality of the solutions, and the trend of contributions in CL in medicine. Several surveys have been conducted on CL using natural images (van de Ven et al., 2022; Wang et al., 2023; Zhou et al., 2023). However, little attention has been paid to the field related to CL and its applications in medicine, a broad and important research area. Moreover, as the trend increases, as shown in Figure 2, a detailed review of this field becomes necessary. Consequently, a survey that captures all augmentations in the field is essential.
1.2 Study Selection
@@
Our selection process involved an extensive search targeting studies centered on the application of CL in medical settings. We initially identified 2100 studies by searching for continual learning and medicine. Subsequently, we refined our focus to 177 relevant computer science and engineering papers from 2018 onwards. These papers were then assessed based on the inclusion and exclusion criteria, where we evaluated their alignment with our research objectives. The inclusion criteria were as follows: (a) addressed both CL and Medical Imaging; (b) proposed solutions for clinical challenges in the medical field; and (c) offered insights or remedies concerning issues related to catastrophic forgetting. Conversely, the exclusion criteria eliminated papers that (a) solely focused on CL without relevance to medical imaging, (b) primarily concentrated on FDA-related topics, and (c) addressed only catastrophic forgetting without relevance to medical imaging or clinical problem solving. Following the study selection process, we curated the final selection of 67 papers. Our final list of studies includes papers that not only explored essential aspects of continual learning in medical imaging but also offered significant insights and solutions for clinical challenges while addressing issues related to catastrophic forgetting.
1.3 Paper Organisation
The rest of this paper is organized as follows. We start by mathematically formalizing CL in Section 2. Then, we discuss and formulate the three types of CL: Task Incremental Learning (Section 2.1), Domain Incremental Learning (Section 2.2), and Class Incremental Learning (Section 2.3). Within this section, we first discuss the theory of these three types of CL before mathematically formulating them with an example. In Section 3, we review the recent advances in CL methods in the medical field and divide the studies on the approach they took to alleviate forgetting. Then, in Section 4, we discussed the overview of the studies along with the current research trend (section 4.1) and the practicality of these studies (section 4.2). Section 5 provides insight into future directions in the field and section 6 concludes the paper.
@@ 1.3 Paper Organisation
2 Background
In this section, we present a comprehensive conceptual framework that includes the mathematical formulation of three primary types of CL: task, domain, and class incremental learning. The framework highlights the potential of these three types and how they can be applied to various advanced medical tasks. The purpose of this framework is to underline the essential concepts of CL, and how diagnostic and prognostic models with CL capabilities are robust and adaptable. The framework showcases the applications of CL in the medical field through Figure 3, thereby highlighting CL as a foundational feature in models that need to incrementally learn from new data sources while retaining previously learned knowledge.
@@ 2 Background
2.1 Task-Incremental Learning
@@
2.2 Domain-Incremental Learning
@@
2.3 Class-Incremental Learning
@@
$\arg\min_{\theta_{\zeta+1}}\left[\mathcal{L}\left(f\left(X_{C_{i}},\theta_{\zeta+1}\right),Y_{C_{i}}\right)+\lambda\Omega\left(\theta_{\zeta+1},\theta_{\zeta}\right)+\beta\Psi\left(\theta_{\zeta+1}\right)\right]$
@@
3 Taxonomy
@@
3.1 Regularization Approaches
Figure 4: Regularization Based Approach: The model is updated by constraining the parameter changes. Assuming $\theta_{new}$ and $\theta_{old}$ are the parameters of the updated and previous models, respectively. The objective is to find an optimal point ($\bullet$) between the parameter space of these models.
@@ 3.1 Regularization Approaches
Regularization-based continual learning offers an effective solution for learning new tasks while preventing catastrophic forgetting of previous ones. The key concept behind this approach is to minimize the drift in the learned feature or parameter space of the model, as shown in Figure 4. By doing so, the model can learn new tasks without inducing significant changes in the parameters that are useful for old tasks, enabling it to retain prior knowledge (Wang et al., 2023). Regularization-based CL incorporates additional loss terms to prevent model drift between old and new tasks. This is typically achieved by maintaining a copy of the model from the previous task or by preserving the model predictions. In medical applications, regularization-based CL are valuable because they do not require access to stored samples, which is a challenge because of privacy concerns. In addition, it does not dynamically increase network size. This makes regularization-based CL models lightweight and versatile, making them particularly suitable for various medical settings. Regularization-based methods fall into two primary categories: parameter regularization and functional regularization, depending on whether they focus on the network parameters or output.
3.1.1 Parameter Regularization
@@
Parameter Regularization focuses on the change in network parameters when learning new tasks. This is achieved by introducing an additional loss term to penalize alterations in the parameters deemed important for the previous tasks. Typically, this involves maintaining a frozen copy of the old model to calculate the change in network parameters with the introduction of new tasks. Several studies have demonstrated the effectiveness of this approach in the medical field. Baweja et al. (2018) incorporated a quadratic penalty in the loss function to address parameter changes important for prior tasks. It follows the Elastic Weight Consolidation (EWC) method using the Fisher information matrix to quantify each parameter’s contribution (Kirkpatrick et al., 2017). Their work comprehensively analyzed the effectiveness of the EWC method in medical imaging applications. Similarly, van Garderen et al. (2019) applied a strategy of penalizing changes in important parameters for brain MRI segmentation, demonstrating its efficacy in transfer learning by pre-training on a high-quality dataset before fine-tuning it on a smaller, lower-quality dataset. This provides an insight into the application of regularization-based CL methods in transfer learning for low-quality datasets, which are common in the medical domain. Another study by Lenga et al. (2020) used EWC and Learning Without Forgetting (LWF) methods for multidomain chest X-ray classification and compared their performance against joint training. Their work showed that the LWF outperformed the EWC approach and achieved a backward transfer on par with joint training on the new domain conducted with 60-80% of the data reply of the previous domain. Chen and Tang (2022) combined the hybrid-averaging operation and EWC into CNN models for breast cancer image classification tasks and found that the ResNet (He et al., 2016) outperforms AlexNet (Krizhevsky et al., 2012) and DenseNet (Huang et al., 2017) in terms of average accuracy and average forgetting. In addition, Derakhshani et al. (2022) and Quarta et al. (2022) offered a benchmark for various regularization-based CL methods, including EWC (Kirkpatrick et al., 2017), LwF (Li and Hoiem, 2017), iCaRL (Rebuffi et al., 2017), MAS (Aljundi et al., 2018), and EEIL (Castro et al., 2018), and showed that iCaRL and EEIL provide superior performance compared to other methods.
@@
More recently, Özgün et al. (2020) proposed a modified version of the memory-aware synapses (MAS) (Aljundi et al., 2018) method by imposing learning rate (LR) constraints on each parameter, instead of direct parameter regularization. In this approach, each parameter or an entire CNN kernel is assigned a specific LR based on its importance for the previous tasks. The authors also proposed freezing the important parameters and fine-tuning the unimportant parameters using LR constraints. Furthermore, Chen et al. (2021) introduced the Targeted Gradient Descent (TGD) method to fine-tune a model without forgetting previous knowledge by leveraging the Kernel Sparsity and Entropy (KSE) metric in a positron emission tomography (PET) reconstruction study. TGD provides an effective method of reusing redundant kernels in a pre-trained CNN for learning new tasks while preserving the balance between the rigidity and plasticity of the models. Another study by Bayasi et al. (2021) used a pruning criterion that allows a fixed network to learn new data domains sequentially over time by detecting the culprit units associated with the wrong classification. These network units are dedicated to the subsequent learning of new domains. MUSCLE (Liao et al., 2022) proposed the addition of model parameter constraints for pretraining a robust backbone in the classification, segmentation, and detection of X-ray images. It utilizes momentum updates between model parameters to prevent catastrophic forgetting during pretraining when moving from one task to another. Ranem et al. (2022) combined a ViT (Dosovitskiy et al., 2020) and nnUNet (Isensee et al., 2021) models for continual hippocampus segmentation using the EWC and PLOP methods(Douillard et al., 2021). Their proposed model utilizes the self-attention mechanism of transformers to maintain knowledge and consequently alleviate catastrophic forgetting, in contrast to the purely convolutional UNet. Nevertheless, interestingly, their experiments showed that applying regularization to the transformer component over time has an adverse effect on knowledge retention, as regularization disrupts the self-attention mechanism. In contrast, it is preferable to discourage substantial changes in the convolutional layers to retain prior knowledge without compromising the plasticity. Furthermore, Zhang et al. (2021, 2023a) proposed a comprehensive importance (CI) method for selecting important parameters based on their contribution to the shape and certainty of output segmentation. They demonstrated that maintaining such parameters can preserve the important shape and semantic information of the segmentation task in CL settings. In another study by Shu et al. (2022), regularization in the gradient space of weights was implemented. This involved updating the model weights using gradients orthogonal to the important parameters for the previous task(s). To further enhance the performance, they incorporated data replay into their methods. Furthermore, Wang and Luo (2022) introduced a technique to mitigate catastrophic forgetting in lung nodule detection studies. Their approach was based on a modified EWC and incorporated feature distillation for enhanced performance. Conversely, Liu et al. (2023) used a frozen copy of the model along with other techniques to adapt an off-the-shelf segmentation model for different anatomical regions. The two model copies were combined by taking a weighted average of the corresponding layers at the end of each training round, allowing for the retention of previous knowledge while acquiring new task-specific information.
3.1.2 Functional Regularization

### v2

# Continual Learning in Medical Imaging: A Survey and Practical Analysis

@@ Abstract
Index Terms:
Continual Learning, Medical, Survey, Types of Conntinal Learning
I Introduction
@@
Deep Learning (DL) algorithms are rapidly gaining relevance in medical imaging, enabling segmentation [1, 2], classification [3], and detection [4] of anatomical structures and anomalies [5, 6] relevant for diagnosis, prediction, or prognosis. In some cases, their capabilities surpass those of human experts [7, 8], making them a central tool in the advancement of the use of medical data to support clinical decisions [9, 10]. Although there are several DL works in medicine, the task is often oversimplified by assuming a static, centralized data source and neglecting the dynamic nature of the evolving data.
In general, a model is trained and tested on different splits of the same dataset, and the models are expected to generalize well to other datasets of the same nature. However, this does not reflect the reality of real-world deployment because of frequent changes in data distribution across various institutions, diagnostic devices, and population demographics [11]. For Instance, in 2020, Google developed a DL system trained on a vast heterogeneous dataset that achieved a high performance in an offline setting for detecting diabetic eye diseases [12]. However, when this system was deployed, its performance drops drastically, highlighting the severity of the problem and emphasizing the need for robust systems. These issues can be partially addressed through domain adaptation and generalization.
@@
Continual learning (CL) has emerged as a promising approach for training deep neural networks in order to overcome these challenges. CL offers a solution that enables neural networks to learn continuously, resulting in a more resilient model capable of accommodating new tasks without forgetting previous tasks. Although the model remains static when deployed in a static environment, it continuously updates itself in a CL setting, thereby yielding a more robust and versatile model. Addressing the challenges mentioned above requires consideration of the rigidity and plasticity of the neural networks. Rigidity, the ability of a neural network to preserve previous knowledge, and plasticity, the ability to learn new tasks are crucial factors. A model with high plasticity can learn new tasks over time, achieving good performance on future tasks but deteriorating performance on older tasks, a phenomenon known as ”catastrophic forgetting”. An ideal neural network should balance rigidity and plasticity to demonstrate its ability to learn new tasks without forgetting prior knowledge. CL has emerged as a promising approach for balancing these properties, enabling neural networks to learn continuously and adapt to evolving data distribution.
I-A Motivation
@@
Significant efforts have been made to develop CL techniques for the natural and medical imaging domains [13, 14, 15, 16, 17]. However, numerous unique challenges arise when dealing with medical data. As illustrated in Figure 1, medical imaging often encounters the challenge of changes in data distribution over time due to both unexpected factors and significant changes in tasks or modalities. Unexpected factors include (a) the incorporation of data from new medical centers, which might vary in acquisition equipment, image reconstruction algorithms, population demographics, and disease expression, and (b) changes within the same site, such as equipment degradation or replacement. Significant changes, on the other hand, include (c) the integration of new modality streams that could enhance downstream tasks and (d) the addition of new tasks, which may require different approaches. Several studies have been conducted to address these challenges. In this work, we aim to conduct a survey of the existing literature and research that may highlight the solutions, the practicality of the solutions, and the trend of contributions in CL in medicine. Several surveys have been conducted on CL using natural images [18, 19, 20]. However, little attention has been paid to the field related to CL and its applications in medicine, a broad and important research area. Moreover, as the trend increases, as shown in Figure 2, a detailed review of this field becomes necessary. Consequently, a survey that captures all aspects in the field is essential.
I-B Study Selection
@@
Our selection process involved an extensive search targeting studies centered on the application of CL in medical settings. We initially identified 2100 studies by searching for continual learning and medicine from various databases and digital libraries including Scopus, Springer and ScienceDirect. Subsequently, we refined our focus to 177 relevant computer science and engineering papers from 2018 onwards. These papers were then assessed based on the inclusion and exclusion criteria, where we evaluated their alignment with our research objectives. The inclusion criteria were as follows: (a) addressed both CL and Medical Imaging; (b) proposed solutions for clinical challenges in the medical field; and (c) offered insights or remedies concerning issues related to catastrophic forgetting. Conversely, the exclusion criteria eliminated papers that (a) solely focused on CL without relevance to medical imaging, (b) primarily concentrated on FDA-related topics, and (c) addressed only catastrophic forgetting without relevance to medical imaging or clinical problem solving. Following the study selection process, we curated the final selection of 67 papers. Our final list of studies includes papers that not only explored essential aspects of continual learning in medical imaging but also offered significant insights and solutions for clinical challenges while addressing issues related to catastrophic forgetting.
@@
I-C Paper Organisation
The rest of this paper is organized as follows. We start by mathematically formalizing CL in Section II. Then, we discuss and formulate the three types of CL: Task Incremental Learning (Section II-A1), Domain Incremental Learning (Section II-A2), and Class Incremental Learning (Section II-A3). Within this section, we first discuss the theory of these three types of CL before mathematically formulating them with an example. In Section III, we review the recent advances in CL methods in the medical field and divide the studies on the approach they took to alleviate forgetting. Then, in Section IV, we discussed the overview of the studies along with the current research trend (section IV-A) and the practicality of these studies (section IV-B). Section V provides insight into future directions in the field and section VI concludes the paper.
II Background
In this section, we present a comprehensive conceptual framework that includes the mathematical formulation of three primary types of CL: task, domain, and class incremental learning. The framework highlights the potential of these three types and how they can be applied to various advanced medical tasks. The purpose of this framework is to underline the essential concepts of CL, and how diagnostic and prognostic models with CL capabilities are robust and adaptable. The framework showcases the applications of CL in the medical field through Figure 3, thereby highlighting CL as a foundational feature in models that need to incrementally learn from new data sources while retaining previously learned knowledge. Then we summarize common evaluation metrics used for continual learning methods.
II-A Types of Incremental Learning
@@
II-A1 Task-Incremental Learning
@@
II-A2 Domain-Incremental Learning
@@
II-A3 Class-Incremental Learning
@@
$\arg\min_{\theta_{\zeta+1}}\left[\mathcal{L}\left(f\left(X_{C_{i}},\theta_{\zeta+1}\right),Y_{C_{i}}\right)+\lambda\Omega\left(\theta_{\zeta+1},\theta_{\zeta}\right)+\beta\Psi\left(\theta_{\zeta+1}\right)\right]$ (4)
@@
II-B Evaluation Metrics
Several evaluation metrics are commonly used in continual learning. These metrics help assess performance across tasks in an incremental learning setting. Below, we summarize some of the most frequently used metrics:
Let $A_{j}^{i}$ (where $i,j=1,2,\dots,T$) represent the accuracy of a model trained after task $j$ and tested on the data from task $i$. Here, $A_{j}^{i}$ refers to the accuracy of the model trained after task $j$ and tested on all previously seen tasks (test data from tasks $0,1,\dots,j$).
The arrows $\uparrow$ and $\downarrow$ denote whether ”higher is better” or ”lower is better,” respectively.
  • Average Accuracy (AvgACC): $\uparrow$
    $\text{AvgACC}=\frac{1}{T}\sum_{i=1}^{T}A^{i}$
  • Final Average Accuracy (FAA): $\uparrow$
    $\text{FAA}=\frac{1}{T}\sum_{i=1}^{T}A_{T}^{i}$
  • Backward Transfer (BWT): $\uparrow$
    $\text{BWT}=\frac{1}{T-1}\sum_{i=2}^{T}\sum_{j=1}^{i-1}\left(A_{j}^{i}-A_{i}^{i}\right)$
  • Average Forgetting (AvgF): $\downarrow$
    $\text{AvgF}=\frac{1}{T-1}\sum_{i=1}^{T-1}\left(\max_{k=1,\dots,T-1}A_{i}^{k}-A_{T}^{i}\right)$
III Taxonomy
Figure 4: Taxonomy of Continual Learning Approaches.
@@ III Taxonomy
III-A Regularization Approaches
@@
Regularization-based continual learning offers an effective solution for learning new tasks while preventing catastrophic forgetting of previous ones. The key concept behind this approach is to minimize the drift in the learned feature or parameter space of the model, as shown in Figure 5. By doing so, the model can learn new tasks without inducing significant changes in the parameters that are useful for old tasks, enabling it to retain prior knowledge [19]. Regularization-based CL incorporates additional loss terms to prevent model drift between old and new tasks. This is typically achieved by maintaining a copy of the model from the previous task or by preserving the model predictions. In medical applications, regularization-based CL are valuable because they do not require access to stored samples, which is a challenge because of privacy concerns. In addition, it does not dynamically increase network size. This makes regularization-based CL models lightweight and versatile, making them particularly suitable for various medical settings. Regularization-based methods fall into two primary categories: parameter regularization and functional regularization, depending on whether they focus on the network parameters or output.
III-A1 Parameter Regularization
@@
Parameter Regularization focuses on the change in network parameters when learning new tasks. This is achieved by introducing an additional loss term to penalize alterations in the parameters deemed important for the previous tasks. Typically, this involves maintaining a frozen copy of the old model to calculate the change in network parameters with the introduction of new tasks. Several studies have demonstrated the effectiveness of this approach in the medical field. [21] incorporated a quadratic penalty in the loss function to address parameter changes important for prior tasks. It follows the Elastic Weight Consolidation (EWC) method using the Fisher information matrix to quantify each parameter’s contribution [22]. Their work comprehensively analyzed the effectiveness of the EWC method in medical imaging applications. Similarly, [23] applied a strategy of penalizing changes in important parameters for brain MRI segmentation, demonstrating its efficacy in transfer learning by pre-training on a high-quality dataset before fine-tuning it on a smaller, lower-quality dataset. This provides an insight into the application of regularization-based CL methods in transfer learning for low-quality datasets, which are common in the medical domain. Another study by [24] used EWC and Learning Without Forgetting (LWF) methods for multidomain chest X-ray classification and compared their performance against joint training. Their work showed that the LWF outperformed the EWC approach and achieved a backward transfer on par with joint training on the new domain conducted with 60-80% of the data replay of the previous domain. [25] combined the hybrid-averaging operation and EWC into CNN models for breast cancer image classification tasks and found that the ResNet [26] outperforms AlexNet [27] and DenseNet [28] in terms of average accuracy and average forgetting. In addition, [29] and [30] offered a benchmark for various regularization-based CL methods, including EWC [22], LwF [31], iCaRL [32], MAS [33], and EEIL [34], and showed that iCaRL and EEIL provide superior performance compared to other methods.
@@
More recently, [35] proposed a modified version of the memory-aware synapses (MAS) [33] method by imposing learning rate (LR) constraints on each parameter, instead of direct parameter regularization. In this approach, each parameter or an entire CNN kernel is assigned a specific LR based on its importance for the previous tasks. The authors also proposed freezing the important parameters and fine-tuning the unimportant parameters using LR constraints. Furthermore, [36] introduced the Targeted Gradient Descent (TGD) method to fine-tune a model without forgetting previous knowledge by leveraging the Kernel Sparsity and Entropy (KSE) metric in a positron emission tomography (PET) reconstruction study. TGD provides an effective method of reusing redundant kernels in a pre-trained CNN for learning new tasks while preserving the balance between the rigidity and plasticity of the models. Another study by [37] used a pruning criterion that allows a fixed network to learn new data domains sequentially over time by detecting the culprit units associated with the wrong classification. These network units are dedicated to the subsequent learning of new domains. MUSCLE [38] proposed the addition of model parameter constraints for pretraining a robust backbone in the classification, segmentation, and detection of X-ray images. It utilizes momentum updates between model parameters to prevent catastrophic forgetting during pretraining when moving from one task to another. [39] combined a ViT [40] and nnUNet [41] models for continual hippocampus segmentation using the EWC and PLOP methods[42]. Their proposed model utilizes the self-attention mechanism of transformers to maintain knowledge and consequently alleviate catastrophic forgetting, in contrast to the purely convolutional UNet. Nevertheless, interestingly, their experiments showed that applying regularization to the transformer component over time has an adverse effect on knowledge retention, as regularization disrupts the self-attention mechanism. In contrast, it is preferable to discourage substantial changes in the convolutional layers to retain prior knowledge without compromising the plasticity. Furthermore, [43, 44] proposed a comprehensive importance (CI) method for selecting important parameters based on their contribution to the shape and certainty of output segmentation. They demonstrated that maintaining such parameters can preserve the important shape and semantic information of the segmentation task in CL settings. In another study by [45], regularization in the gradient space of weights was implemented. This involved updating the model weights using gradients orthogonal to the important parameters for the previous task(s). To further enhance the performance, they incorporated data replay into their methods. Furthermore, [46] introduced a technique to mitigate catastrophic forgetting in lung nodule detection studies. Their approach was based on a modified EWC and incorporated feature distillation for enhanced performance. Conversely, [47] used a frozen copy of the model along with other techniques to adapt an off-the-shelf segmentation model for different anatomical regions. The two model copies were combined by taking a weighted average of the corresponding layers at the end of each training round, allowing for the retention of previous knowledge while acquiring new task-specific information.
Figure 5: Regularization Based Approach: The model is updated by constraining the parameter changes. Assuming $\theta_{new}$ and $\theta_{old}$ are the parameters of the updated and previous models, respectively. The objective is to find an optimal point between the parameter space of these models.
III-A2 Functional Regularization